OpenStudy (anonymous):

What is the value of a in the following equation: 7.05a - 2= -4.25a + 26.25? 2.1 10.1 2.5 5.2

OpenStudy (haseeb96):

7.05a + 4.25a =26.25+2 11.3a=28.25 a=28.25\11.3 a=2.5 so C is the correct answer

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@Haseeb96 Thanks :) I like how you show your work, it helps me :)

OpenStudy (anonymous):

So I can learn it myself :)